OpenStudy (anonymous):

Can anyone please tell me two numbers that, when added give you 0, and when multiplied, give you a positive number. Thanks!

OpenStudy (dan815):

x+y=0 x*y>0 x=-y -y*y>0 -y^2>0 y=complex number

OpenStudy (anonymous):

Sorry but I really didn't get that. Could you please explain it more simply? I just need 2 numbers

OpenStudy (dan815):